OpenStudy (anonymous):

Which ordered pair is a solution of the equation 4x - 2y = 16?

OpenStudy (anonymous):

There are many ordered pairs. Do you have any to chose from?

OpenStudy (anonymous):

a. (-5, -3) b. (-5, 3) c. (6, -4) d. (6, 4)

OpenStudy (meowlover17):

Okay plug those in for x, and y and see which one makes the equation true.

OpenStudy (anonymous):

Take each one and plug them in to see if it is true. I will do a a. (-5, -3) 4x - 2y = 16 4(-5) - 2(-3) = 16 -20 - 2(-3) = 16 -20 - 6 = 16 -25=16 This is not true so it is not our answer. Can you do the rest to find the answer?